OpenStudy (nikki_skellington):

Jamie has a deck of 60 sports cards, of which some are baseball cards and some are football cards. Jamie pulls out a card randomly from the deck, records its type, and replaces it in the deck. Jamie has already recorded three baseball cards and nine football cards. Based on these data, what is, most likely, the number of baseball cards in the deck? 12 15 24 30

OpenStudy (nikki_skellington):

is 24 correct?

OpenStudy (math2vcap):

30

OpenStudy (nikki_skellington):

um...explanation plz

OpenStudy (nikki_skellington):

@agent0smith could u help?

OpenStudy (math2vcap):

it has to be a 50 50 chance

OpenStudy (nikki_skellington):

could u tell me how u got 30

OpenStudy (math2vcap):

60/2

OpenStudy (math2vcap):

is 30

OpenStudy (nikki_skellington):

...thats not how u do it....

OpenStudy (nikki_skellington):

but thx for trying :/

OpenStudy (agent0smith):

Jamie has already recorded three baseball cards and nine football cards. prob of baseball = 3/12 = 1/4 Multiply that by the total number of cards

OpenStudy (nikki_skellington):

?

OpenStudy (math2vcap):

its 15 because for every 12 cards 3 are baseball cards and 12 goes into 60 5 times

OpenStudy (nikki_skellington):

are u sooo sure its 15 cuz i rlly need to get this right

OpenStudy (math2vcap):

yes

OpenStudy (math2vcap):

am I right?

OpenStudy (nikki_skellington):

o w8 its is 15!

OpenStudy (nikki_skellington):

thx guys!

OpenStudy (agent0smith):

" Jamie has already recorded three baseball cards and nine football cards. prob of baseball = 3/12 = 1/4 Multiply that by the total number of cards " What do you need help understanding here?